### v1

# Magnetic field dynamics in presence of Hall conductivity, and thermodiffusion

Abstract
This anisotropy of kinetic coefficients in presence of a magnetic field is represented by so-called Hall currents, which appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions. In many papers the Hall currents had been considered in different approximate approaches. We derive equations, describing dynamics of the magnetic field in presence of Hall currents, using a standard electrodynamic consideration. We consider collisional media, and take into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component. The influence of the Hall currents on the magnetic field structure and damping is considered in simple models. In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery”.
1 Introduction
In presence of a magnetic field kinetic coefficients, describing charged particles, became anisotropic. It is related to the plasma, and to the metals. This anisotropy represents appearance of so-called Hall currents. They appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions. There is an ample literature where different aspects of the Hall currents were investigated, in particular, their influence on the magnetic field structure and damping. The most striking feature of the Hall currents is their property of preserving entropy, and absence of a heat production. Therefore the magnetic field damping due to Hall currents may happen only indirectly, by changing the magnetic field structure. In many papers (see i.g. [1],[2], [3],[4],[5],[6]) the account if the Hall current had been done with some approximations, including works with consideration of GR effects [7].
Here we derive equations, using exact electrodynamic consideration, describing dynamics of the magnetic field in presence of Hall currents. We consider collisional media, like fully ionized plasma in the space and in the laboratory. We have taken into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component.
In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery” [8, 9]. The influence of the Hall currents on the magnetic field structure and damping is considered in simple models.
1.1 MHD equations in presence of Hall conductivity
The MHD equations include the Maxwell equations without a displacement current [10], and a generalized Ohm law for anisotropic conductivity in presence of the magnetic field in the form:
$\frac{1}{c}\frac{\partial\bf B}{\partial t}=-{\bf\nabla\,\times\,E},\qquad{\bf\nabla\cdot B}=0,$ (1)
${\bf j}=\frac{c}{4\pi}{\bf\nabla\,\times\,B},$ (2)
${\bf j}=\sigma_{ij}\{{\bf E}+\frac{1}{c}{\bf[v\times B]}\}.$ (3)
Anisotropic conductivity tensor $\sigma_{ij}$ has a relatively simple form in the first Chapman-Enskog approximation [11]. Finally the expression in (3) may be written in the form [3]-[6]:
${\bf j}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr{]}.$ (4)
For a non-relativistic, non-degenerate, fully ionized plasma we have the following scalar coefficients of its conductivity $\sigma_{0}$, in the absence of the magnetic field, see [11],[16],[21]:
$\sigma_{0}=\frac{32e^{2}n_{e}}{3\pi m_{e}}\tau_{e},\,\,{\rm with}\,\,\tau_{e}=\frac{3}{4}\sqrt{\frac{m_{e}}{2\pi}}\frac{(kT)^{3/2}}{Z^{2}e^{4}n_{N}\Lambda}.$ (5)
Here $\omega=(eB/m_{e}c)$ is the electron Larmor frequency, $\tau_{e}$ is a time between electron-ion collision, in Lorentz approximation; $n_{e}$, $n_{N}$ are concentrations of the electrons and nuclei with atomic number Z, $\Lambda$ is a Coulomb logarithm, ${\bf E}$, ${\bf E^{\prime}}$ are electric fields in the laboratory, and comoving coordinate system, respectively.
${\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]}.$ (6)
In Cartesian coordinates $(x,y,z)$ with
${\bf E^{\prime}}=(E^{\prime}_{x},\,E^{\prime}_{y},\,E^{\prime}_{z});\qquad{\bf B}=(B_{x},\,B_{y},\,B_{z}),$ (7)
the generalized Ohm law (4) is represented by the following three equations:
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{x}=E^{\prime}_{x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{x}-$
$\frac{\omega\tau_{e}}{B}(B_{y}E^{\prime}_{z}-B_{z}E^{\prime}_{y}),$ (8)
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{y}=E^{\prime}_{y}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{y}-$
$\frac{\omega\tau_{e}}{B}(B_{z}E^{\prime}x-B_{x}E^{\prime}_{z}),$ (9)
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{z}=E^{\prime}_{z}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{z}-$
$\frac{\omega\tau_{e}}{B}(B_{x}E^{\prime}_{y}-B_{y}E^{\prime}_{x}).$ (10)
For deriving a dynamic equation for the magnetic field ${\bf B}$ evolution in conditions of anisotropic conductivity we need to solve the system of the above linear equations for finding a dependence ${\bf E^{\prime}=E^{\prime}(j,B)}$, for substituting it in the first equation (1). We have found the solution in the following form:
$E^{\prime}_{x}=\frac{1}{\sigma_{0}}[j_{x}-\frac{\omega\tau_{e}}{B}(B_{z}j_{y}-B_{y}j_{z})],\qquad\qquad$ (11)
$E^{\prime}_{y}=\frac{1}{\sigma_{0}}[j_{y}-\frac{\omega\tau_{e}}{B}(B_{x}j_{z}-B_{z}j_{x})],\qquad\qquad$ (12)
$E^{\prime}_{z}=\frac{1}{\sigma_{0}}[j_{z}-\frac{\omega\tau_{e}}{B}(B_{y}j_{x}-B_{x}j_{y})].\qquad\qquad$ (13)
With account of (2), this solution is written in the vector form as:
${\bf E^{\prime}}=\frac{\bf j}{\sigma_{0}}-\frac{\omega\tau_{e}}{B\sigma_{0}}{\bf[B\times j]}=\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$ (14)
This equation had been written in [10],[18], where a third-party E.D.S. was also included.
Substituting (14) into (1), with account of (6), we obtain the equation for the evolution of a magnetic field at anisotropic electric conductivity due to Hall current, in the form [2],[3],[19],[5],[6]:
${\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}+\frac{c^{2}\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf\nabla}\times[{\bf B\times[\nabla\times B]}]}.$ (15)
The hydrodynamic part of MHD equations [20] (continuity, motion energy, and heat propagation) should be added to the Eqs.(15), at Ohm’s law (4), for solving problems appearing in modeling of events in the space and in a laboratory.
The equation (15) is describing the evolution of the magnetic field in the medium moving with center mass velocity $\bf v$ relative to the laboratory frame. In presence of a magnetic field $\bf B$ it has an anisotropic conductivity, with the Hall component represented by the last term in (15).
It is interesting to compare (15) with the equation (39) from the paper [1], where the first term to the left has the same dependence on $\bf B$, as the last term in (15), but has a different coefficient. The authors of [1] have called it as ”Hall drift”. Actually this term is connected with Lorentz transformation of the electrical field E from the laboratory frame to the co-moving one [10]. The Hall effect is connected with the anisotropic electrical conductivity in the conducting medium in presence of the magnetic field, which is included in the generalized Ohms law [11], given here in Eq.(4). In [1] the equation (39) is applied to the metallic static body with the electric current $\bf j$, which is obtained from the Lorentz transformation equation by identifying ${\bf j}$= $e{\bf v}$.
This procedure is not correct, because the center mass ${\bf v}$, connected with the moving frame is actually zero, and the electric current should be found from the Ohms law, like in Eq.(4). The model in [1] is almost the same as in the case of the plasmic body, because the electrical current is produced mainly by the electron motion in the medium moving with the velocity ${\bf v}$, which is zero in this case.
The difference between plasmic and metallic body consist of difference in the scalar electrical conductivity $\sigma_{0}$, and in different ways of supplying a static state. For the metal bar the motion is described by the mechanic law, and a plasmic liquid body is governed by the MHD equations. According to (15), in the ideal plasma with infinite $\sigma_{0}$, there is no damping of the magnetic field, no influence of the Hall currents on its motion and in the co-moving frame the electric field is zero. In [1] the so-called ”Hall drift” term remains non-zero in the ideal plasma, what leads to artificial behavior of the magnetic field in it.
2 Equations in the presence of thermodiffusion, and external electro-moving force (EMF)
With account of thermodiffusion and external electrical field ${\bf E^{{}^{\prime\prime}}}$, the Ohm’s law has the following form, when using the same approximation as in the derivation of (4), and remaining only the term with a temperature gradient ${\bf\nabla}T$ in the diffusion vector ${\bf d}$, see [15]:
${\bf j}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}}\biggr{]}$
$\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}}\biggr{]}$ (16)
$+\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr{]}.$
Here $\lambda_{0}$ is a scalar thermodiffusion coefficient, which for a non-relativistic, non-degenerate plasma in absence of a magnetic field, in Lorenz approximation is written as
$\lambda_{0}=en_{e}\mu^{(1)}=\frac{16ken_{e}}{\pi m_{e}}\tau_{e}.$ (17)
The linear equations, connecting the components of the electrical current $j_{i}$, electrical field in the co-moving frame $E^{\prime}_{i}$ and temperature gradient $\frac{\partial T}{\partial x_{i}}$ are written as ($i=x,y,z$):
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{x}=E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}[B_{x}(E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x})+B_{y}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})\qquad$ (18)
$+B_{z}(E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z})]B_{x}-\frac{\omega\tau_{e}}{B}[B_{y}(E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z})-B_{z}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})]\qquad$
$+\frac{\lambda_{0}}{\sigma_{0}}\biggl{[}\frac{\partial T}{\partial x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}\frac{\partial T}{\partial x}+B_{y}\frac{\partial T}{\partial y}+B_{z}\frac{\partial T}{\partial z})B_{x}-\frac{\omega\tau_{e}}{B}(B_{y}\frac{\partial T}{\partial z}-B_{z}\frac{\partial T}{\partial y})\biggr{]},$
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{y}=E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}[B_{x}(E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x})+B_{y}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})\qquad$ (19)
$+B_{z}(E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z})]B_{y}-\frac{\omega\tau_{e}}{B}[B_{z}(E^{\prime}x+E^{{}^{\prime\prime}}_{x})-B_{x}(E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z})]\qquad$
$+\frac{\lambda_{0}}{\sigma_{0}}\biggl{[}\frac{\partial T}{\partial y}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}\frac{\partial T}{\partial x}+B_{y}\frac{\partial T}{\partial y}+B_{z}\frac{\partial T}{\partial z})B_{y}-\frac{\omega\tau_{e}}{B}(B_{z}\frac{\partial T}{\partial x}-B_{x}\frac{\partial T}{\partial z})\biggr{]},$
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{z}=E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}[B_{x}(E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x})+B_{y}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})\qquad$ (20)
$+B_{z}(E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z})]B_{z}-\frac{\omega\tau_{e}}{B}[B_{x}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})-B_{y}E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x})]\qquad$
$+\frac{\lambda_{0}}{\sigma_{0}}\biggl{[}\frac{\partial T}{\partial z}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}\frac{\partial T}{\partial x}+B_{y}\frac{\partial T}{\partial y}+B_{z}\frac{\partial T}{\partial z})B_{z}-\frac{\omega\tau_{e}}{B}(B_{x}\frac{\partial T}{\partial y}-B_{y}\frac{\partial T}{\partial x})\biggr{]}.$
Introducing
$\tilde{E}_{x}=E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x}+\frac{\lambda_{0}}{\sigma_{0}}\frac{\partial T}{\partial x},\quad\tilde{E}_{y}=E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y}+\frac{\lambda_{0}}{\sigma_{0}}\frac{\partial T}{\partial y},\quad\tilde{E}_{z}=E^{\prime}_{z}+E^{{}^{\prime\prime}}_{z}+\frac{\lambda_{0}}{\sigma_{0}}\frac{\partial T}{\partial z},$ (21)
we obtain the solution of the system (18)-(20) in the form:
$\tilde{E}_{x}=\frac{1}{\sigma_{0}}[j_{x}-\frac{\omega\tau_{e}}{B}(B_{z}j_{y}-B_{y}j_{z})],\qquad\qquad$ (22)
$\tilde{E}_{y}=\frac{1}{\sigma_{0}}[j_{y}-\frac{\omega\tau_{e}}{B}(B_{x}j_{z}-B_{z}j_{x})],\qquad\qquad$ (23)
$\tilde{E}_{z}=\frac{1}{\sigma_{0}}[j_{z}-\frac{\omega\tau_{e}}{B}(B_{y}j_{x}-B_{x}j_{y})].\qquad\qquad$ (24)
With account of (2), this solution is written in the vector form as:
${\bf\tilde{E}}={\bf E^{\prime}+E^{{}^{\prime\prime}}}+\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T=\frac{\bf j}{\sigma_{0}}-\frac{\omega\tau_{e}}{B\sigma_{0}}{\bf[B\times j]}$ (25)
$=\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$ (26)
In presence of thermodiffusion and external electrical field $\bf E^{{}^{\prime\prime}}$ we obtain, after applying the curl procedure to the equation (26), and using (1), the following equation:
$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+c{\bf\nabla\times E^{{}^{\prime\prime}}}+c\nabla\biggl{(}\frac{\lambda_{0}}{\sigma_{0}}\biggr{)}\times{\nabla T}\qquad$ (27)
$+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}-\frac{c^{2}}{4\pi}[{\bf\nabla}(\frac{1}{\sigma_{0}})\times{\bf\nabla\times B}]+\frac{c^{2}}{4\pi}{\bf\nabla}\times\big{[}\frac{\omega\tau_{e}}{\sigma_{0}B}{\bf B}\times[\nabla\times{\bf B}]\big{]}.$
The Eq.(27) is valid for variable values of the parameters $\lambda_{0}$ and $\sigma_{0}$. At constant value $\lambda_{0}/\sigma_{0}=3k/2e$ in the fully ionized Lorenz plasma, the Eq.(27) does not contain the temperature terms. It means that in this conditions thermodiffusion is not participating directly in the magnetic field evolution. In this case the Eq.(27) coincides with the similar equation from [10].
At variable transport coefficients the temperature term in the Eq.(27) may be non-zero even at initially zero values of ${\bf B}$ and ${\bf E}$, what leads to creation of a seed magnetic field in the non-magnetized media. At the function $\lambda_{0}/\sigma_{0}=f(\rho,T)$, the seed magnetic field is created at non-zero value of
$\frac{\partial f}{\partial\rho}[\nabla\rho\times\nabla T]$
It is representing the equation for Biermann battery [8, 9], which is derived here by the solution of Boltzmann kinetic equation.
In the system with temperature gradients, external electric field ${\bf E^{{}^{\prime\prime}}}$ and external magnetic field ${\bf B_{ex}}$, connected with external electric current ${\bf j_{ex}}$, there are internal fields ${\bf E_{in}}$, ${\bf B_{in}}$ and internal electric current ${\bf j_{in}}$. The internal fields are connected by the first Maxwell equation (1). Finally, we obtain the following system of equation for determination of these values:
$\frac{1}{c}\frac{{\partial\bf B}_{in}}{\partial t}=-{\bf\nabla\,\times\,E},\quad{\bf\nabla\cdot B}_{in}=0,\quad{\bf j_{in}}=\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{in},\quad\qquad\qquad$ (28)
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]},\qquad$ (29)
${\bf j}_{in}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr{]},\quad{\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]},\quad$ (30)
$\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{ex}^{v}={\bf j_{ex}},\quad{\bf\nabla\cdot B}_{ex}=0,\qquad{\bf B}_{ex}={\bf B}_{ex}^{v}+{\bf B}_{ex}^{vf},\qquad\qquad$ (31)
${\bf j}_{ex}=\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr{]}\qquad\qquad\qquad$ (32)
$+\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}]}\biggr{]},\qquad\qquad\qquad$
${\bf j}={\bf j_{in}}+{\bf j_{ex}},\qquad{\bf B}={\bf B}_{in}+{\bf B}_{ex}.\qquad\qquad\qquad\qquad$ (33)
Here ${\bf E}$, ${\bf E^{\prime}}$ are inner electric fields measured in the laboratory, and comoving coordinate system, respectively. In addition there are given external electrical field ${\bf E^{{}^{\prime\prime}}}$, external current ${\bf j_{ex}}$, creating a vortex external magnetic field ${\bf{B_{ex}^{v}}}$; and independent vortex-free magnetic field ${\bf{B_{ex}^{vf}}}$.
There is a complicate situation due to presence of a external fields, currents and thermodiffusion. All currents are producing magnetic fields, which, together with the external field ${\bf{B_{ex}^{vf}}}$, are taking part in the formation of Hall currents. The currents are produced in different processes, but the magnetic field which they have created, is acting together, as one vector ${\bf B}$. These properties are represented in the system of equations (28)-(33).
3 Simple models with an action of the Hall current in the presence of thermodiffusion
3.1 Magnetized plasma cylinder
Here we obtain from (28) - (33) the equations, describing the behavior of the plasma cylinder (see Fig. 1) with a uniform magnetic field $B_{0}$ along its axis, and a radial temperature gradient. In the case of cylinder symmetry $\frac{\partial}{\partial z}=\frac{\partial}{\partial\phi}$, the only non-zero parameters of heat flux and electric current left are: $q_{r},q_{\phi},j_{r},j_{\phi}$.
Figure 1: Conducting cylinder with Hall current $j_{Hall}$, depending on the magnitude of the radial temperature gradient, and external constant magnetic field $B_{0}$ along its axis. The induced magnetic field $B_{1}$ is determined by the Hall current. $R_{1}$ is the radius of the central heated region with constant temperature $T_{0}$. Toroidal region, colored in gray, contains Hall current and associated magnetic field, which has an opposite direction to the external field $B_{0}$, decreasing the resulting field along the cylinder.
In absence of the external electric field $E^{{}^{\prime\prime}}$ and mutually perpendicular vectors ${\bf B,\nabla}T$, equation (32) is written as follows:
${\bf j_{ex}}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}({{\bf\nabla}T}-\frac{\omega\tau_{e}}{B}[{\bf B\times\nabla}T]).$ (34)
The components of the electrical current density $j_{ex}$ in a cylinder with $B_{z}$ and radial temperature gradient are written as:
$j_{ex,r}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr},\qquad j_{ex,\phi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (35)
From eq. (31) we have the equation for the inner magnetic field, induced by the Hall current, as:
$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\lambda_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (36)
Solution of this equation at different input parameters was obtained in [21]. The magnetic field $B_{z}^{v}$, induced by the Hall current due to thermodiffusion, in all cases had a sign, opposite to the external field $B_{z,ex}^{v}$.
When the external radial electric field $E_{r,ex}$ is present in the cylinder, instead of the radial temperature gradient, then instead of the Eq.(36), we have the following equation for the induced field $B_{z}^{v}$:
$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\sigma_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}E_{r,ex},\quad\sigma_{0}=\frac{32}{3\pi}\frac{e^{2}n_{e}\tau_{e}}{m_{e}}.$ (37)
Here $\sigma_{e}$ is the scalar plasma electrical conductivity. This field has the same opposite sign in relation to the external field $B_{z,ex}^{vf}$.
3.2 Magnetized plasma torus
Figure 2: Torus with a initial electric field $E_{0}$, that produce circular magnetic field $B_{0}$. $B_{0}$ and temperature gradient $\nabla T$ create Hall electric current $j_{Hall}$ in opposite direction to the $E_{0}$. The induced magnetic field $B_{1}$ is determined by the Hall current $j_{Hall}$.
Let us consider a circular plasma torus, in the center of which there is a linear electrical current with a linear density ${J_{z}}$, directed perpendicular to the symmetry plain of the torus $(x,y)$, see Fig.2. The cross section of the torus, which includes $z$-axis, is supposed to be circular, with a radius $R_{t}$. In the symmetry plane $z=0$ the torus is situated between the radius $R_{out}$ and $R_{in}$, where:
$R_{out}=R_{in}+2R_{t}.$
Using Biot–Savart law, the magnetic field of the infinite linear current is represented by concentric circles around the current in the plane perpendicular to it. This external field has only $\phi$ component and is written as [22]:
$B_{\phi}^{ex}=\frac{2J_{z}}{cr}.$ (38)
Torus surface is defined by the 4-th order equation. It is formed by a rotation of a circle with the radius $R_{t}$, around a $z$-axis, when the center of the circle has the radius $R=R_{in}+R_{t}$. The equation of this surface is written as follows
$(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=R_{t}^{2},\quad{\rm which\,\,reduced\,\,to}\quad$
$(x^{2}+y^{2}+z^{2}+R^{2}-R_{t}^{2})^{2}=4R^{2}(x^{2}+y^{2}).\,\,\,$ (39)
By heating the central ring of the torus and formation of the temperature gradient ${\bf\nabla T}$, we produce a heat flux inside the torus directed to the surface of the torus in the plane, containing the $z$ axis. In the cylinder coordinate system $r,\phi,z$ the non-zero temperature gradient takes place along the coordinate $\rho$, defined as:
$\rho^{2}=(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=(r-R)^{2}+z^{2},\quad 0\leq\rho\leq R_{t}.$ (40)
With this definition we have  $\rho d\rho=(r-R)dr+zdz$,
$\frac{\partial T}{\partial r}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial r}=\frac{r-R}{\rho}\frac{dT}{d\rho},\qquad\frac{\partial T}{\partial z}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial z}=\frac{z}{\rho}\frac{dT}{d\rho}.$ (41)
For constant heat flux $Q$ through the unit length of the torus, the temperature gradient along the small torus radius $\rho$ is determined as
$q=\frac{Q}{2\pi}=\frac{\kappa_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\quad\kappa_{0}=\frac{320}{3\pi}\frac{k^{2}Tn_{e}\tau_{e}}{m_{e}}.$ (42)
Here the scalar heat conductivity coefficient $\kappa_{0}$ is determined in [17]. The thermodiffusion is forming an electrical current along the temperature gradient over the $\rho$ coordinate $j_{ex,\rho}$, and a hall current $j_{ex,\psi}$ along the small circle of the torus:
$j_{ex,\rho}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\qquad j_{ex,\psi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho}.$ (43)
Due to the axial symmetry, the magnetic field in the torus is produced only by its Hall component $j_{ex,\psi}$, namely, by the circular current along the small circle of the torus, see Fig.2. The torus with such current is used in the construction of the devices for obtaining controlled thermonuclear reaction [23], where such currents are concentrated in wire around the torus, creating the toroidal field going around the torus. In our model similar currents of Hall origin are created in the plasma of the torus itself, when the current is created in the wire circles, with the current $I$, along the torus, with $n$ wire circles on the unit length. The total circular current around the small torus radius is equal to $2\pi RnI$. The magnetic field circulation along a big circle on the distance $r$ from the central big circle is equal to $2\pi rB$. From the circulation theorem we obtain then the poloidal field inside such torus as [24]:
$B_{\phi}(r)=nI\frac{R}{rc},\quad R_{in}<r<R.$ (44)
In the plasma torus there is a circular Hall current density $j_{ex,\psi}$, defined in (43). The magnetic field formed by the Hall current could be considered as sum of thin circular currents, similar to the wire current in the laboratory torus, considered above. The total circular current on the radial distance $r_{0}$ from the center is equal to ($j_{ex,\psi}2\pi r_{0}d\rho$). The magnetic field circulation from this current along the torus at the radius $r<r_{0}$ is equal to $2\pi rB$. So the magnetic field $dB_{\phi}$ produced by this sheet of current at the radius $r$ is written as:
$dB_{\phi}(r)=j_{ex,\psi}\frac{r_{0}}{rc}d\rho,\quad R_{in}<r<r_{0}.$ (45)
It follows from (44), after substituting ($nI$) by ($j_{ex,\psi}(\rho)d\rho$), and $R$ by $r_{0}$. The Hall magnetic field from the whole current in the plasma torus is obtained by integration of (45) over the torus thickness as
$B_{\phi}(r)=\int_{0}^{R-R_{in}}j_{ex,\psi}\frac{r_{0}}{rc}d\rho,=\frac{1}{rc}\int_{0}^{R-R_{in}}j_{ex,\psi}r_{0}d\rho,\qquad R_{in}<r<R.$ (46)
On the symmetry plane $z=0$ there is a connection $\rho=R-r_{0}$, $d\rho=-dr_{0}$. For constant current density we have after integration:
[…]
It is easy to show that the magnetic field produced by the Hall current has a sign opposite to the field from the central current (Fig.2). Field decreasing is minimal on the central line, and its maximal value is in the surface of the torus.
When the external radial electric field $E_{\rho,ex}$ is present in the torus instead of the temperature gradient, then instead of expression in the Eq.(43), we should use the following one:
$j_{ex,\rho}=\frac{\sigma_{0}E_{\rho,ex}}{1+(\omega\tau_{e})^{2}},\qquad j_{ex,\psi}=-\sigma_{0}\frac{\omega\tau_{e}E_{\rho,ex}}{1+(\omega\tau_{e})^{2}}.$ (48)
4 Discussion
Equations describing magnetic field dynamics in fully ionized nonuniform plasma are derived rigorously, with account of Hall currents and thermodiffusion effects. It the system of equation written here, the internally and externally produced values of magnetic and electrical fields ${\bf B}$ and ${\bf E}$, and also electrical current ${\bf j}$ are distinctly separated. In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, that is modeling the action of the mechanism, known as ”Biermann battery”. Application of these equations is done for examples of plasma cylinder and plasma torus. In both cases the externally induced electric current may be formed by the temperature gradient (thermodiffusion), or by external electric field ( battery or accumulator). In all cases the internal magnetic field, produced by the Hall currents, has a direction opposite to the externally induced magnetic field.
The derived equations could be used for theoretical modeling of magnetic field behavior in astrophysical objects, like different types of stars, fully ionized galactic gas, etc. Another application may be connected with numerical modeling of laboratory experiments with production and acceleration of magnetized, high energy, non-uniform plasma, some of which are related to laboratory astrophysics.
Acknowledgments
This work was partially supported by RSCF 23-12-00198 grant.

### v2

@@
-Magnetic field dynamics in presence of Hall conductivity, and thermodiffusion
+Magnetic field dynamics in presence of thermo-diffusion and Hall conductivity
 Abstract
-This anisotropy of kinetic coefficients in presence of a magnetic field is represented by so-called Hall currents, which appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions.
-In many papers the Hall currents had been considered in different approximate approaches.
-We derive equations, describing dynamics of the magnetic field in presence of Hall currents, using a standard electrodynamic consideration.
-We consider collisional media, and take into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component.
-The influence of the Hall currents on the magnetic field structure and damping is considered in simple models.
-In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery”.
+Anisotropy of kinetic coefficients in presence of a magnetic field is represented by Hall currents, which appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions.
+We derive equations, describing dynamics of the magnetic field in presence of thermo-diffusion with Hall currents, using a standard electrodynamic consideration.
+The influence of the Hall currents, at presence of thermo-diffusion, on the magnetic field structure is considered in simple models.
+The equation is derived, which includes additional term for a seed magnetic field creation in the non-magnetized plasma, due to thermo-diffusion.
+This equation describes the action of the seed magnetic field creation by the mechanism known as ”Biermann battery”.
 1 Introduction
-In presence of a magnetic field kinetic coefficients, describing charged particles, became anisotropic.
-It is related to the plasma, and to the metals.
-This anisotropy represents appearance of so-called Hall currents.
-They appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions.
-There is an ample literature where different aspects of the Hall currents were investigated, in particular, their influence on the magnetic field structure and damping.
-The most striking feature of the Hall currents is their property of preserving entropy, and absence of a heat production.
-Therefore the magnetic field damping due to Hall currents may happen only indirectly, by changing the magnetic field structure.
-In many papers (see i.g.
-[1],[2], [3],[4],[5],[6]) the account if the Hall current had been done with some approximations, including works with consideration of GR effects [7].
-Here we derive equations, using exact electrodynamic consideration, describing dynamics of the magnetic field in presence of Hall currents.
-We consider collisional media, like fully ionized plasma in the space and in the laboratory.
-We have taken into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component.
-In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery” [8, 9].
-The influence of the Hall currents on the magnetic field structure and damping is considered in simple models.
-1.1 MHD equations in presence of Hall conductivity
+In presence of a magnetic field, kinetic coefficients, describing charged particles, became anisotropic.
+It is related to plasma, as well as to metals.
+Such anisotropy represents appearance of Hall currents, which appear in a collisional plasma, due to action of the Lorentz force on the charged particles between collisions.
+The characteristic feature of Hall currents is their property of preserving entropy, and absence of a heat production.
+Therefore the influence of Hall currents on a magnetic field damping may happen only indirectly, by changing the magnetic field structure.
+Transport coefficients in the non-degenerate magnetized plasma had been found in [11], [12], using solution of the Boltzmann kinetic equation by Chapmen-Enskog method [10].
+Thermo-diffusion and Hall currents had been taken into account in these papers, together with analysis of different physical effects.
+The equation for the excitation of a seed magnetic field in the non-magnetized plasma, obtained there, is connected with a pressure gradient term in the diffusion vector $\bf d_{e}$} [10], [12]
+${\bf d_{e}}=\frac{e}{kT}\large[{\bf E}+\frac{1}{c}[{\bf v}\times{\bf B}]\large]+\frac{{\bf\nabla}P_{e}}{P_{e}}=\frac{e}{kT}\large[{\bf E}+\frac{1}{c}[{\bf v}\times{\bf B}]+\frac{{\bf\nabla}P_{e}}{en_{e}}\large]$ (1)
+We have derived here the equation for the excitation of a seed magnetic field in presence of thermo-diffusion, and have obtained the additional term in the equation for seed field excitation.
+The creation of a seed magnetic field is a very important astrophysical problem.
+The presence of magnetic field everywhere in the universe is explained by different mechanisms of its amplification, but these processes could happen only after creation of initial seed field in the initially non-magnetized universe,
+Development of a laboratory astrophysics is aimed to model astrophysical events in a terrestrial laboratory.
+MHD simulations with Hall current had been performed in [37], to simulate the laboratory experiments.
+Numerical calculations of the magnetic field evolution in neutron star crust had been done in [4]-[5] in presence of the Hall effect.
+Here similar equations are derived in presence of a temperature gradient, external action, and thermo-diffusion.
+We have separated magnetic and electric fields on internal and external one, depending on their origin: internal electrical currents and electrical charge distribution, and external source.
+Those could be the electrical field produced by battery, accumulators etc., and magnetic field originated by external currents, or external non-vortex magnetic field.
+In presence of thermo-diffusion, the additional channel for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery” [7, 8].
+Simple models are considered for illustration of Hall current’s influence on magnetic filed structure, including the action of thermo-diffusion.
+2 MHD equations in presence of Hall conductivity
 The MHD equations include the Maxwell equations without a displacement current [9], and a generalized Ohm law for anisotropic conductivity in presence of the magnetic field in the form:
@@
 ${\bf j}=\frac{c}{4\pi}{\bf\nabla\,\times\,B},$ (3)
-${\bf j}=\sigma_{ij}\{{\bf E}+\frac{1}{c}{\bf[v\times B]}\}.$ (3)
-Anisotropic conductivity tensor $\sigma_{ij}$ has a relatively simple form in the first Chapman-Enskog approximation [11]. Finally the expression in (3) may be written in the form [3]-[6]:
+${\bf j}=\sigma_{ij}{{\bf E^{\prime}}}.$ (4)
+Anisotropic conductivity tensor $\sigma_{ij}$ has a relatively simple form in the first Chapman-Enskog approximation, $\bf E^{\prime}$} is the the electrical field in the co-moving coordinate system.
+The expression in (4) may be written in the form [3]-[6]:
 ${\bf j}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr].$ (5)
@@
 $\sigma_{0}=\frac{32e^{2}n_{e}}{3\pi m_{e}}\tau_{e},\,\,{\rm with}\,\,\tau_{e}=\frac{3}{4}\sqrt{\frac{m_{e}}{2\pi}}\frac{(kT)^{3/2}}{Z^{2}e^{4}n_{N}\Lambda}.$ (6)
-Here $\omega=(eB/m_{e}c)$ is the electron Larmor frequency, $\tau_{e}$ is a time between electron-ion collision, in Lorentz approximation; $n_{e}$, $n_{N}$ are concentrations of the electrons and nuclei with atomic number Z, $\Lambda$ is a Coulomb logarithm, ${\bf E}$, ${\bf E^{\prime}}$ are electric fields in the laboratory, and comoving coordinate system, respectively.
-${\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]}.$ (6)
-In Cartesian coordinates $(x,y,z)$ with
-${\bf E^{\prime}}=(E^{\prime}_{x},\,E^{\prime}_{y},\,E^{\prime}_{z});\qquad{\bf B}=(B_{x},\,B_{y},\,B_{z}),$ (7)
-the generalized Ohm law (4) is represented by the following three equations:
-$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{x}=E^{\prime}_{x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{x}-$
-$\frac{\omega\tau_{e}}{B}(B_{y}E^{\prime}_{z}-B_{z}E^{\prime}_{y}),$ (8)
-$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{y}=E^{\prime}_{y}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{y}-$
-$\frac{\omega\tau_{e}}{B}(B_{z}E^{\prime}x-B_{x}E^{\prime}_{z}),$ (9)
-$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{z}=E^{\prime}_{z}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{z}-$
-$\frac{\omega\tau_{e}}{B}(B_{x}E^{\prime}_{y}-B_{y}E^{\prime}_{x}).$ (10)
-For deriving a dynamic equation for the magnetic field ${\bf B}$ evolution in conditions of anisotropic conductivity we need to solve the system of the above linear equations for finding a dependence ${\bf E^{\prime}=E^{\prime}(j,B)}$, for substituting it in the first equation (1).
-We have found the solution in the following form:
-$E^{\prime}_{x}=\frac{1}{\sigma_{0}}[j_{x}-\frac{\omega\tau_{e}}{B}(B_{z}j_{y}-B_{y}j_{z})],\qquad\qquad$ (11)
-$E^{\prime}_{y}=\frac{1}{\sigma_{0}}[j_{y}-\frac{\omega\tau_{e}}{B}(B_{x}j_{z}-B_{z}j_{x})],\qquad\qquad$ (12)
-$E^{\prime}_{z}=\frac{1}{\sigma_{0}}[j_{z}-\frac{\omega\tau_{e}}{B}(B_{y}j_{x}-B_{x}j_{y})].\qquad\qquad$ (13)
-With account of (2), this solution is written in the vector form as:
+Here $\omega=(eB/m_{e}c)$ is the electron Larmor frequency, $\tau_{e}$ is a time between electron-ion collision, in Lorentz approximation; $n_{e}$, $n_{N}$ are concentrations of the electrons and nuclei with atomic number Z, $\Lambda$ is a Coulomb logarithm, $\bf E$}, $\bf E^{\prime}$} are electric fields in the laboratory, and comoving coordinate system, with account of the pressure gradient term, respectively .
+The comoving system is moving with velocity $\bf v$} relative to the laboratory system, together with matter, so that [9],[12]
+${\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]}+\frac{{\bf\nabla}P_{e}}{en_{e}}.$ (7)
+From equations (2)-(5) follows the expression for $\bf E^{\prime}$}, as a function of $\bf j,\,B$}, with account of Hall currents, in the form [9],[17]
 ${\bf E^{\prime}}=\frac{\bf j}{\sigma_{0}}-\frac{\omega\tau_{e}}{B\sigma_{0}}{\bf[B\times j]}=\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$ (8)
-This equation had been written in [10],[18], where a third-party E.D.S. was also included.
-Substituting (14) into (1), with account of (6), we obtain the equation for the evolution of a magnetic field at anisotropic electric conductivity due to Hall current, in the form [2],[3],[19],[5],[6]:
-${\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}+\frac{c^{2}\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf\nabla}\times[{\bf B\times[\nabla\times B]}]}.$ (15)
+Substituting (8) into (2), with account of (7), leads to the equation for the evolution of a magnetic field [12], [2],[3],[18],[5],[6], at anisotropic electric conductivity due to Hall current, as
+$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}$ (9)
+$+\frac{c^{2}\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf\nabla}\times[{\bf B\times[\nabla\times B]}]-\frac{ck}{en_{e}}[{\bf\nabla}n_{e}\times{\bf\nabla}T].$
 The hydrodynamic part of MHD equations [19] (continuity, motion energy, and heat propagation) should be added to the Eqs.(9), at Ohm’s law (5), for solving problems appearing in modeling of events in the space and in a laboratory.
 The equation (9) is describing the evolution of the magnetic field in the medium moving with center mass velocity $\bf v$ relative to the laboratory frame.
-In presence of a magnetic field $\bf B$ it has an anisotropic conductivity, with the Hall component represented by the last term in (15).
-It is interesting to compare (15) with the equation (39) from the paper [1], where the first term to the left has the same dependence on $\bf B$, as the last term in (15), but has a different coefficient.
+It is interesting to compare (9) with the equation (39) from the paper [1], where the first term to the right has the same dependence on $\bf B$, as the second term from the end in (9), but has a different coefficient.
 The authors of [1] have called it as ”Hall drift”.
@@
 In [1] the equation (39) is applied to the metallic static body with the electric current $\bf j$, which is obtained from the Lorentz transformation equation by identifying $\bf j$}= $e{\bf v}$.
-This procedure is not correct, because the center mass ${\bf v}$, connected with the moving frame is actually zero, and the electric current should be found from the Ohms law, like in Eq.(4).
-The model in [1] is almost the same as in the case of the plasmic body, because the electrical current is produced mainly by the electron motion in the medium moving with the velocity ${\bf v}$, which is zero in this case.
-The difference between plasmic and metallic body consist of difference in the scalar electrical conductivity $\sigma_{0}$, and in different ways of supplying a static state.
-For the metal bar the motion is described by the mechanic law, and a plasmic liquid body is governed by the MHD equations.
-According to (15), in the ideal plasma with infinite $\sigma_{0}$, there is no damping of the magnetic field, no influence of the Hall currents on its motion and in the co-moving frame the electric field is zero.
+This procedure is not correct, because the center mass $\bf v$}, related to the moving frame, is actually zero.
+According to (9), in the ideal plasma with infinite $\sigma_{0}$, and absence of the pressure gradient term, there is no damping of the magnetic field, no influence of the Hall currents on its behavior, and in the co-moving frame the electric field is zero.
 In [1] the so-called ”Hall drift” term remains non-zero in the ideal plasma, what leads to artificial behavior of the magnetic field in it.
-2 Equations in the presence of thermodiffusion, and external electro-moving force (EMF)
-With account of thermodiffusion and external electrical field ${\bf E^{{}^{\prime\prime}}}$, the Ohm’s law has the following form, when using the same approximation as in the derivation of (4), and remaining only the term with a temperature gradient ${\bf\nabla}T$ in the diffusion vector ${\bf d}$, see [15]:
+3 Equations in the presence of thermo-diffusion, and external electro-moving force (EMF)
+With account of thermo-diffusion and external electrical field $\bf E^{{}^{\prime\prime}}$}, see [9], [10],[11],[12], [13], the Ohm’s law has the following form:
 ${\bf j}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}}\biggr]$
@@
 $\lambda_{0}=en_{e}\mu^{(1)}=\frac{16ken_{e}}{\pi m_{e}}\tau_{e}.$ (11)
-The linear equations, connecting the components of the electrical current $j_{i}$, electrical field in the co-moving frame $E^{\prime}_{i}$ and temperature gradient $\frac{\partial T}{\partial x_{i}}$ are written as ($i=x,y,z$):
+The linear equations, connecting the components of the electrical current $j_{i}$, electrical field in the co-moving frame $E^{\prime}_{i}$, external electrical field $E^{\prime\prime}_{i}$, and temperature gradient $\frac{\partial T}{\partial x_{i}}$ are written as ($i=x,y,z$):
 $\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{x}=E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}[B_{x}(E^{\prime}_{x}+E^{{}^{\prime\prime}}_{x})+B_{y}(E^{\prime}_{y}+E^{{}^{\prime\prime}}_{y})\qquad$ (12)
@@
 $\tilde{E}_{z}=\frac{1}{\sigma_{0}}[j_{z}-\frac{\omega\tau_{e}}{B}(B_{y}j_{x}-B_{x}j_{y})].\qquad\qquad$ (18)
-With account of (2), this solution is written in the vector form as:
+With account of (3), (7), this solution is written in the vector form as:
 ${\bf\tilde{E}}={\bf E^{\prime}+E^{{}^{\prime\prime}}}+\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T=\frac{\bf j}{\sigma_{0}}-\frac{\omega\tau_{e}}{B\sigma_{0}}{\bf[B\times j]}$ (19)
 $=\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$
-${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$ (26)
-In presence of thermodiffusion and external electrical field $\bf E^{{}^{\prime\prime}}$ we obtain, after applying the curl procedure to the equation (26), and using (1), the following equation:
-$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+c{\bf\nabla\times E^{{}^{\prime\prime}}}+c\nabla\biggl{(}\frac{\lambda_{0}}{\sigma_{0}}\biggr{)}\times{\nabla T}\qquad$ (27)
+${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}$ (20)
+$-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}-\frac{{\bf\nabla}P_{e}}{en_{e}}.$
+In presence of thermodiffusion and external electrical field $\bf E^{{}^{\prime\prime}}$ we obtain, after applying the curl procedure to the equation (20), and using (2), the following equation at $P_{e}=n_{e}kT$:
+$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+c{\bf\nabla\times E^{{}^{\prime\prime}}}+c\nabla\biggl(\frac{\lambda_{0}}{\sigma_{0}}\biggr)\times{\nabla T}-\frac{ck}{en_{e}}[{\bf\nabla}n_{e}\times{\bf\nabla}T]$ (21)
 $+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}-\frac{c^{2}}{4\pi}[{\bf\nabla}(\frac{1}{\sigma_{0}})\times{\bf\nabla\times B}]+\frac{c^{2}}{4\pi}{\bf\nabla}\times\big[\frac{\omega\tau_{e}}{\sigma_{0}B}{\bf B}\times[\nabla\times{\bf B}]\big].$
+3.1 The Biermann battery
+Two last terms in the upper line of Eq.(21) may create a seed magnetic field in the non- magnetized media.
+The first term is related to the thermo-diffusion, and the second one is connected with a current produces by a pressure gradient (barodiffusion).
+The seed magnetic field is created when the surfaces of the constant concentration $n_{e}$ don’t coincide with surfaces of constant temperature,
+$\nabla n_{e}\times\nabla T\neq 0.$
+These two terms are representing the Biermann battery [7, 8].
 The Eq.(21) is valid for variable values of the parameters $\lambda_{0}$ and $\sigma_{0}$.
-At constant value $\lambda_{0}/\sigma_{0}=3k/2e$ in the fully ionized Lorenz plasma, the Eq.(27) does not contain the temperature terms.
-It means that in this conditions thermodiffusion is not participating directly in the magnetic field evolution.
-In this case the Eq.(27) coincides with the similar equation from [10].
-At variable transport coefficients the temperature term in the Eq.(27) may be non-zero even at initially zero values of ${\bf B}$ and ${\bf E}$, what leads to creation of a seed magnetic field in the non-magnetized media.
-At the function $\lambda_{0}/\sigma_{0}=f(\rho,T)$, the seed magnetic field is created at non-zero value of
+At constant value $\lambda_{0}/\sigma_{0}=3k/2e=const$, in the fully ionized non-degenerate Lorenz plasma, the first temperature term is zero.
+At in this conditions thermo-diffusion is not participating in creation of the seed magnetic field.
+When this term is variable,
+$\lambda_{0}/\sigma_{0}=f(\rho,T)$
+the seed magnetic field is created by a thermo-diffusion at non-zero value of
 $\frac{\partial f}{\partial\rho}[\nabla\rho\times\nabla T]$
-It is representing the equation for Biermann battery [8, 9], which is derived here by the solution of Boltzmann kinetic equation.
+In neutron star crust the electrons are strongly degenerated, and the conditions of the seed field creation are changed.
+Let us consider the case of strongly degenerate, non-relativistic electrons. In this case the coefficients of conductivity and thermo-diffusion, calculated in [20], have the form
+$\sigma_{0}^{d}=\frac{e^{2}n_{e}}{m_{e}}\tau_{d},\qquad\lambda_{0}^{d}=\frac{4\pi^{3}e}{(9\pi)^{1/3}}\frac{n_{e}^{1/3}k^{2}T}{h^{2}}\tau_{d}.$ (22)
+Here
+$\tau_{d}=\frac{3}{32\pi^{2}}\frac{h^{3}n_{e}}{Z^{2}e^{4}m_{e}n_{N}\Lambda},\qquad h\,\,\,\,{\mbox{ is the Planck constant}}.$ (23)
+The ratio of transport coefficients is written as
+$\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}=\frac{k}{e}\,\frac{4\pi^{3}}{(9\pi)^{1/3}}\frac{kTm_{e}}{h^{2}n_{e}^{2/3}}$ (24)
+what may give a non-zero value $\nabla n_{e}\times\nabla T\neq 0.$
+For strongly degenerate plasma the baro-diffusion term in (21) should be taken, using (7), with the equation of state in the form (see [20])
+$P_{e}=2kT\big(\frac{2kTm_{e}}{h^{2}}\big)^{3/2}G_{5/2}(x_{0}),\quad n_{e}=2\big(\frac{2kTm_{e}}{h^{2}}\big)^{3/2}G_{3/2}(x_{0}).$ (25)
+We obtain from [21],[16]
+$x_{0}=\frac{\varepsilon_{fe}}{kT}=\frac{(3\pi^{2}n_{e})^{2/3}h^{2}}{8\pi^{2}m_{e}kT}\gg 1,\qquad\qquad$ (26)
+$G_{5/2}(x_{0})\approx\frac{x_{0}^{5/2}}{\Gamma(7/2)}+\frac{\pi^{2}}{4}\frac{x_{0}^{1/2}}{\Gamma(5/2)},\quad G_{3/2}(x_{0})\approx\frac{x_{0}^{3/2}}{\Gamma(5/2)}+\frac{\pi^{2}}{12}\,\frac{x_{0}^{-1/2}}{\Gamma(3/2)},$
+$P_{e}=\frac{(3\pi^{2})^{2/3}}{20\pi^{2}}\frac{h^{2}n_{e}^{5/3}}{m_{e}}\big(1+\frac{5\pi^{2}}{8x_{0}^{2}}\big)=P_{ed}\big(1+\frac{5\pi^{2}}{8x_{0}^{2}}\big).$ (27)
+The creation of the seed magnetic field in the strongly degenerate plasma is described by the equation
+$\frac{\partial{\bf B}}{\partial t}=c\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)\times{\nabla T}+\frac{c}{e}\,\nabla\times\biggl(\frac{{\bf\nabla}P_{e}}{n_{e}}\biggr)$ (28)
+We obtain from (24) and (27)
+$\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)=\frac{k}{e}\,\frac{4\pi^{3}}{(9\pi)^{1/3}}\frac{km_{e}}{h^{2}}\biggl(\frac{\nabla T}{n_{e}^{2/3}}-\frac{2}{3}\frac{T\nabla n_{e}}{n_{e}^{5/3}}\biggr),$ (30)
+$\nabla\times\biggl(\frac{{\bf\nabla}P_{e}}{n_{e}}\biggr)\,=-\frac{\nabla n_{e}}{n_{e}^{2}}\times{\bf\nabla P_{e}}=-2\frac{P_{ed}}{n_{e}^{2}}\frac{5\pi^{2}}{8x_{0}^{2}T}\nabla n_{e}\times{\nabla T}.$ (31)
+With account of (30),(31) we have from (28)
+$\frac{\partial{\bf B}}{\partial t}=-\biggl[\frac{4\pi^{8/3}}{3^{2/3}}\frac{c}{e}\frac{k^{2}Tm_{e}}{h^{2}\,n_{e}^{5/3}}+\frac{8\,\pi^{7/3}}{3^{5/3}}\,\frac{c}{e}\,\frac{k^{2}Tm_{e}}{h^{2}n_{e}^{5/3}}\biggr](\nabla n_{e}\times{\nabla T})$ (32)
+Using (27), this equation is rewritten as (see also [26])
+$\frac{\partial{\bf B}}{\partial t}=-\biggl[40.7109+18.5312\biggr]\frac{c}{e}\frac{k^{2}Tm_{e}}{h^{2}\,n_{e}^{5/3}}(\nabla n_{e}\times{\nabla T})$ (33)
+In (33) the left term in the brackets is related to baro-diffision, and the right one is for thermo-diffusion. The input of both terms here is comparable, with a constant ratio.
+3.2 The system of MHD equations, related to magnetic field behavior
 In the system with temperature gradients, external electric field $\bf E^{{}^{\prime\prime}}$} and external magnetic field $\bf B_{ex}$}, connected with external electric current $\bf j_{ex}$}, there are internal fields $\bf E_{in}$}, $\bf B_{in}$} and internal electric current $\bf j_{in}$}.
@@
 Finally, we obtain the following system of equation for determination of these values:
-$\frac{1}{c}\frac{{\partial\bf B}_{in}}{\partial t}=-{\bf\nabla\,\times\,E},\quad{\bf\nabla\cdot B}_{in}=0,\quad{\bf j_{in}}=\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{in},\quad\qquad\qquad$ (28)
-${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]},\qquad$ (29)
-${\bf j}_{in}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr{]},\quad{\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]},\quad$ (30)
-$\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{ex}^{v}={\bf j_{ex}},\quad{\bf\nabla\cdot B}_{ex}=0,\qquad{\bf B}_{ex}={\bf B}_{ex}^{v}+{\bf B}_{ex}^{vf},\qquad\qquad$ (31)
-${\bf j}_{ex}=\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr{]}\qquad\qquad\qquad$ (32)
-$+\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}]}\biggr{]},\qquad\qquad\qquad$
+$\frac{1}{c}\frac{{\partial\bf B}_{in}}{\partial t}=-{\bf\nabla\,\times\,E},\quad{\bf\nabla\cdot B}_{in}=0,\quad{\bf j_{in}}=\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{in},\qquad$ (34)
+${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}\qquad$
+$-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}-\frac{{\bf\nabla}P_{e}}{en_{e}},\quad\qquad$ (35)
+${\bf j}_{in}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr],\qquad$ (36)
+${\bf E^{\prime}}={\bf E}+\frac{1}{c}{{\bf[v\times B]}}+\frac{{\nabla}P_{e}}{en_{e}},\,\,\quad\qquad$ (37)
+$\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{ex}^{v}={\bf j_{ex}},\quad{\bf\nabla\cdot B}_{ex}=0,\quad{\bf B}_{ex}={\bf B}_{ex}^{v}+{\bf B}_{ex}^{vf},\qquad$ (38)
+${\bf j}_{ex}=\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr]\quad$ (39)
+$+\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}]}\biggr],\hskip 22.99988pt$
 ${\bf j}={\bf j_{in}}+{\bf j_{ex}},\qquad{\bf B}={\bf B}_{in}+{\bf B}_{ex}.\qquad\qquad\qquad\qquad$ (40)
-Here ${\bf E}$, ${\bf E^{\prime}}$ are inner electric fields measured in the laboratory, and comoving coordinate system, respectively.
+Here $\bf E$}, $\bf E^{\prime}$} are inner electric fields measured in the laboratory, and in the co-moving coordinate system, respectively.
 In addition there are given external electrical field $\bf E^{{}^{\prime\prime}}$}, external current $\bf j_{ex}$}, creating a vortex external magnetic field $\bf{B_{ex}^{v}}$}; and independent vortex-free magnetic field $\bf{B_{ex}^{vf}}$}.
-There is a complicate situation due to presence of a external fields, currents and thermodiffusion.
+There is a complicate situation due to presence of a external fields, currents and thermo-diffusion.
 All currents are producing magnetic fields, which, together with the external field $\bf{B_{ex}^{vf}}$}, are taking part in the formation of Hall currents.
@@
 These properties are represented in the system of equations (34)-(40).
-3 Simple models with an action of the Hall current in the presence of thermodiffusion
+4 Several applications
+4.1 Thrusters
+Thrusters are reactive engines at low power, and very high efficiency, based on the electromagnetic acceleration of plasma.
+A conception of this engine was suggested in the year 1957 by A.I.
+Morozov [22].
+A simplified version of their action is based on the consideration of the ideal plasma behavior inside the mutually perpendicular external electric $\bf E^{{}^{\prime\prime}}$} and magnetic $\bf B_{ex}$} fields, where the internal electrical field in the laboratory system $\bf E$} is zero.
+For the ideal plasma with infinite conductivity $\sigma_{0}$ we have from (3.2) the relation, determining the equilibrium plasma velocity v in these conditions
+${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}=0.$ (41)
+This equation has an exact solution
+${\bf v}=c\frac{\bf E^{{}^{\prime\prime}}\times B}{B^{2}}.$ (42)
+As followed from (34), at zero $\bf E$} the inner magnetic field $\bf B_{in}$}=0 therefore, we take here ${\bf B}={\bf B_{ex}}$.
+The review’s devoted to the 30th anniversary of small plasma thrusters (SPT) operation in space are given in Refs.
+[23] (about the conceptual development of stationary plasma thrusters), and [24] (about stationary plasma thrusters operate in space).
+Usually we have situations when $E\ll B$.
+so that the speed of the out flowing plasma $v\ll c$.
+If, by some reason, the value of $\bf E$} is approaching that of $\bf B$}, the speed of the out flowing plasma is approaching light speed $c$.
+In this situation the approximate Lorentz transformation formulae should be substituted by its exact variant [25], and instead of (41), (42) we obtain following relations for the geometry where, in Cartesian coordinates $(x,y,z)$
+${\bf v}=(V,0,0),\quad{\bf E}=(0,E,0),\quad{\bf B}=(0,0,B),$ (43)
+the following result
+$E=-\frac{V}{c}\frac{B}{\sqrt{1-V^{2}/c^{2}}}-E^{{}^{\prime\prime}}=0.$ (44)
+This equation has an exact solution
+$V={c}\frac{E^{{}^{\prime\prime}}}{\sqrt{B_{ex}^{2}+{E^{{}^{\prime\prime}}}^{2}}}.$ (45)
+4.2 Neutron star crust
+Biermann battery effects connected with a magnetic field behavior, and its possible role in the seed magnetic field creation in the neutron star (NS) crust, have been considered in [26].
+The authors used transport coefficients from the works [27], [28],[29], where numerical estimations of some integral have been made.
+Here we obtain a fully analytic solution of this problem, using some approximation.
+In the middle aged NS crust a temperature is moderate, and a density is very high, so the electrons are strongly degenerate and ultrarelativistic.
+Following [27], we extend approximately the application of the Eq.
+(28) to the relativistic electrons, by using in the transport coefficients, the effective electron mass $m_{*}$, instead of $m_{e}$, with
+$m_{*}\approx{m_{e}}{\sqrt{1+\frac{p_{fe}^{2}}{(m_{e}c)^{2}}}},\qquad p_{fe}=\bigg(\frac{3n_{e}}{8\pi}\biggr)^{1/3}h.$ (46)
+In this case the coefficients of conductivity and thermo-diffusion, with account of (46), are taken in the form
+$\sigma_{0}^{du}=\frac{e^{2}n_{e}}{m_{*}}\tau_{du},\qquad\lambda_{0}^{du}=\frac{4\pi^{3}e}{(9\pi)^{1/3}}\frac{n_{e}^{1/3}k^{2}T}{h^{2}}\tau_{du},$ (47)
+with
+$\tau_{du}=\frac{3}{32\pi^{2}}\frac{h^{3}n_{e}}{Z^{2}e^{4}m_{*}n_{N}\Lambda},\qquad h\,\,\,\,{\mbox{ is the Planck constant}}.$ (48)
+The ratio of transport coefficients is written as
+$\frac{\lambda_{0}^{du}}{\sigma_{0}^{du}}=\frac{k}{e}\,\frac{4\pi^{8/3}}{3^{2/3}}\frac{kTm_{*}}{h^{2}n_{e}^{2/3}},$ (49)
+with possible non-zero value of $\nabla n_{e}\times\nabla T\neq 0.$
+For the baro-diffusion term in (7), (3) of degenerate electrons the equation of state in the relativistic limit is written in the form (see [16])
+$P_{d}=\frac{1}{8}\biggl(\frac{3}{\pi}\biggr)^{1/3}c\,h\,n_{e}^{4/3}\biggl[1+\frac{2\pi^{2/3}}{3^{5/3}}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}n_{e}^{-2/3}\biggr]$ (50)
+The creation of the seed magnetic field in the strongly degenerate plasma is described by the equation (28) We obtain from (49) and (50)
+$\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)=\frac{k}{e}\,\frac{4\pi^{8/3}}{3^{2/3}}\frac{km_{*}}{h^{2}}\biggl(\frac{\nabla T}{n_{e}^{2/3}}-\frac{2}{3}\frac{T\nabla n_{e}}{n_{e}^{5/3}}\biggr),$ (51)
+$\nabla\times\biggl(\frac{\nabla P_{d}}{n_{e}}\biggr)=-\frac{\nabla n_{e}}{n_{e}^{2}}\times{\nabla P_{d}}=-\frac{\pi^{1/3}ch}{2\cdot 3^{4/3}n_{e}^{4/3}T}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}(\nabla n_{e}\times{\nabla T}).$ (52)
+With account of (51),(52) we have from (28)
+$\frac{\partial{\bf B}}{\partial t}=-\biggl[\frac{ck}{e}\,\frac{8\pi^{8/3}}{3^{5/3}}\frac{kTm_{*}}{h^{2}n_{e}^{5/3}}+\frac{c}{e}\,\frac{\pi^{1/3}ch}{2\cdot 3^{4/3}n_{e}^{4/3}T}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}\biggr](\nabla n_{e}\times{\nabla T})$ (53)
+This equation is rewritten as (see also [26])
+$\frac{\partial{\bf B}}{\partial t}=-\frac{ck}{e}\,\frac{2\pi^{7/3}}{3^{4/3}n_{e}^{4/3}}\biggl(\frac{kT}{h\,c}\biggr)\biggl(1+\frac{4\pi^{1/3}}{3^{1/3}}\frac{m_{*}c}{hn_{e}^{1/3}}\biggr)(\nabla n_{e}\times{\nabla T}).$ (54)
+In this consideration the baro-diffusion term is based on exact equation of state for relativistic electrons, while in the thermo-diffusion term we used corrected non-relativistic kinetic coefficients, similar to [27].
+In this situation the physical interpretation of terms in the brackets of (54) is not so clear.
 4.3 Magnetized plasma cylinder
-Here we obtain from (28) - (33) the equations, describing the behavior of the plasma cylinder (see Fig.
-1) with a uniform magnetic field $B_{0}$ along its axis, and a radial temperature gradient.
-In the case of cylinder symmetry $\frac{\partial}{\partial z}=\frac{\partial}{\partial\phi}$, the only non-zero parameters of heat flux and electric current left are: $q_{r},q_{\phi},j_{r},j_{\phi}$.
+The model of the magnetized plasmic cylinder with a radial temperature gradient was investigated in [20].
+The temperature gradient created the thermo-diffusion electrical current, in the external uniform magnetic field along the cylinder axis (see Fig.1).
+The equations describing the structure of the magnetic field in presence of the external thermo-diffusion force, used in [20], follow from Eqs.
+(34) - (40).
+It was obtained in [20], that thermo-diffusional Hall currents in this situation create magnetic field of the sign, opposite the the external one, and can substantially decrease the resulting magnetic field.
+This behavior is in accordance with the Lenz’s law [30] in electrodynamic.
 Figure 1: Conducting cylinder with Hall current $j_{Hall}$, depending on the magnitude of the radial temperature gradient, and external constant magnetic field $B_{0}$ along its axis.
@@
 Toroidal region, colored in gray, contains Hall current and associated magnetic field, which has an opposite direction to the external field $B_{0}$, decreasing the resulting field along the cylinder.
-In absence of the external electric field $E^{{}^{\prime\prime}}$ and mutually perpendicular vectors ${\bf B,\nabla}T$, equation (32) is written as follows:
-${\bf j_{ex}}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}({{\bf\nabla}T}-\frac{\omega\tau_{e}}{B}[{\bf B\times\nabla}T]).$ (34)
-The components of the electrical current density $j_{ex}$ in a cylinder with $B_{z}$ and radial temperature gradient are written as:
-$j_{ex,r}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr},\qquad j_{ex,\phi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (35)
-From eq. (31) we have the equation for the inner magnetic field, induced by the Hall current, as:
-$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\lambda_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (36)
-Solution of this equation at different input parameters was obtained in [21].
-The magnetic field $B_{z}^{v}$, induced by the Hall current due to thermodiffusion, in all cases had a sign, opposite to the external field $B_{z,ex}^{v}$.
-When the external radial electric field $E_{r,ex}$ is present in the cylinder, instead of the radial temperature gradient, then instead of the Eq.(36), we have the following equation for the induced field $B_{z}^{v}$:
-$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\sigma_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}E_{r,ex},\quad\sigma_{0}=\frac{32}{3\pi}\frac{e^{2}n_{e}\tau_{e}}{m_{e}}.$ (37)
-Here $\sigma_{e}$ is the scalar plasma electrical conductivity. This field has the same opposite sign in relation to the external field $B_{z,ex}^{vf}$.
 4.4 Magnetized plasma torus
-Figure 2: Torus with a initial electric field $E_{0}$, that produce circular magnetic field $B_{0}$.
-$B_{0}$ and temperature gradient $\nabla T$ create Hall electric current $j_{Hall}$ in opposite direction to the $E_{0}$.
-The induced magnetic field $B_{1}$ is determined by the Hall current $j_{Hall}$.
-Let us consider a circular plasma torus, in the center of which there is a linear electrical current with a linear density ${J_{z}}$, directed perpendicular to the symmetry plain of the torus $(x,y)$, see Fig.2.
-The cross section of the torus, which includes $z$-axis, is supposed to be circular, with a radius $R_{t}$.
-In the symmetry plane $z=0$ the torus is situated between the radius $R_{out}$ and $R_{in}$, where:
+The plasma torus configurations are used in construction of devices for thermonuclear energy production (Tokamak, Stellarator) [31],[32].
+Bilding of ITER (International Thermonuclear Experimental Reactor) is based on the Tokamak intallation [33].
+In astrophysics the plasma torus may be formed around compact objects (white dwarfs, neutron stars, black holes) during accretion in the binary system.
+The giant magnetized plasma torus configuration, formed around the central nuclei in the active galactic nuclei and quasars, are considered now as a most plausible model, explaining many different observational features [34],[35].
+Here we consider the structure of a circular plasma torus, using Eqs.
+(34) -(40).
+Through the center of the torus there is a linear electrical current, situated perpendicular to the symmetry plain of the torus $(x,y)$, with a linear current $J_{z}^{ex}$}, see Fig.2.
+The cross section of the torus, perpendicular to $z$-axis, is supposed to be circular.
+In the symmetry plane $z=0$ the torus occupies the space between the radius $R_{out}$ and $R_{in}$.
+The cross section of the torus along $z$-axis is also taken as a circular, with the radius $R_{t}$, so that
 $R_{out}=R_{in}+2R_{t}.$
+Figure 2: Torus with a initial electric current $j_{0}$, that produce circular magnetic field $B_{0}$.
+$B_{0}$ and temperature gradient $\nabla T$ create Hall electric current $j_{Hall}$.
+The induced magnetic field $B_{1}$ is determined by the Hall current $j_{Hall}$ and is oriented in the opposite direction to the $B_{0}$.
 Using Biot–Savart law, the magnetic field of the infinite linear current is represented by concentric circles around the current in the plane perpendicular to it.
 This external field has only $\phi$ component and is written as [30]:
-$B_{\phi}^{ex}=\frac{2J_{z}}{cr}.$ (38)
-Torus surface is defined by the 4-th order equation.
+$B_{\phi}^{ex}=\frac{2J_{z}^{ex}}{cr}.$ (55)
+The torus surface is defined by the 4-th order equation.
 It is formed by a rotation of a circle with the radius $R_{t}$, around a $z$-axis, when the center of the circle has the radius $R=R_{in}+R_{t}$.
 The equation of this surface is written as follows
-$(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=R_{t}^{2},\quad{\rm which\,\,reduced\,\,to}\quad$
-$(x^{2}+y^{2}+z^{2}+R^{2}-R_{t}^{2})^{2}=4R^{2}(x^{2}+y^{2}).\,\,\,$ (39)
-By heating the central ring of the torus and formation of the temperature gradient ${\bf\nabla T}$, we produce a heat flux inside the torus directed to the surface of the torus in the plane, containing the $z$ axis.
-In the cylinder coordinate system $r,\phi,z$ the non-zero temperature gradient takes place along the coordinate $\rho$, defined as:
-$\rho^{2}=(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=(r-R)^{2}+z^{2},\quad 0\leq\rho\leq R_{t}.$ (40)
-With this definition we have  $\rho d\rho=(r-R)dr+zdz$,
-$\frac{\partial T}{\partial r}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial r}=\frac{r-R}{\rho}\frac{dT}{d\rho},\qquad\frac{\partial T}{\partial z}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial z}=\frac{z}{\rho}\frac{dT}{d\rho}.$ (41)
-For constant heat flux $Q$ through the unit length of the torus, the temperature gradient along the small torus radius $\rho$ is determined as
-$q=\frac{Q}{2\pi}=\frac{\kappa_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\quad\kappa_{0}=\frac{320}{3\pi}\frac{k^{2}Tn_{e}\tau_{e}}{m_{e}}.$ (42)
+$(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=R_{t}^{2},\quad{\rm which\,\,is\,\,reduced\,\,to}\quad$
+$(x^{2}+y^{2}+z^{2}+R^{2}-R_{t}^{2})^{2}=4R^{2}(x^{2}+y^{2}).\,\,\,\quad$ (56)
+By heating the central ring of the torus and formation of the temperature gradient $\nabla T$}, a heat flux is created inside the torus along its smaller radius $\rho$, directed to the surface of the torus.
+In the cylinder coordinate system $r,\phi,z$ the the coordinate $\rho$ is defined as:
+$\rho^{2}=(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=(r-R)^{2}+z^{2},\quad 0\leq\rho\leq R_{t},$ (57)
+so that in the plane $z=0$,  $\rho d\rho=(r-R)dr$, for $R_{in}<r<R$.
+For a constant heat flux $Q$ through the unit length of the large circle of the torus, the temperature gradient along the small torus radius $\rho$ is determined as
+$q=\frac{Q}{2\pi\rho}=-\frac{\kappa_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\quad\kappa_{0}=\frac{320}{3\pi}\frac{k^{2}Tn_{e}\tau_{e}}{m_{e}}.$ (58)
 Here the scalar heat conductivity coefficient $\kappa_{0}$ is determined in [16].
-The thermodiffusion is forming an electrical current along the temperature gradient over the $\rho$ coordinate $j_{ex,\rho}$, and a hall current $j_{ex,\psi}$ along the small circle of the torus:
-$j_{ex,\rho}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\qquad j_{ex,\psi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho}.$ (43)
-Due to the axial symmetry, the magnetic field in the torus is produced only by its Hall component $j_{ex,\psi}$, namely, by the circular current along the small circle of the torus, see Fig.2.
-The torus with such current is used in the construction of the devices for obtaining controlled thermonuclear reaction [23], where such currents are concentrated in wire around the torus, creating the toroidal field going around the torus.
-In our model similar currents of Hall origin are created in the plasma of the torus itself, when the current is created in the wire circles, with the current $I$, along the torus, with $n$ wire circles on the unit length.
-The total circular current around the small torus radius is equal to $2\pi RnI$.
-The magnetic field circulation along a big circle on the distance $r$ from the central big circle is equal to $2\pi rB$.
-From the circulation theorem we obtain then the poloidal field inside such torus as [24]:
-$B_{\phi}(r)=nI\frac{R}{rc},\quad R_{in}<r<R.$ (44)
-In the plasma torus there is a circular Hall current density $j_{ex,\psi}$, defined in (43).
+The thermo-diffusion is forming the electrical current along the temperature gradient along the $\rho$ coordinate $j_{in,\rho}$, and a hall current $j_{in,\psi}$ along the small circle of the torus:
+$j_{in,\rho}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\qquad j_{in,\psi}=\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho}.$ (59)
+Due to the axial symmetry, the magnetic field inside the torus is produced only by the Hall component $j_{in,\psi}$,corresponding to a circular current along the small circle of the torus, see Fig.2.
+The torus with such current is used in the construction of the devices for obtaining controlled thermonuclear reaction [32], where these currents are concentrated in the wire wound up in the torus around its small radius.
+Such current is creating the toroidal field along the big circle the torus.
+In our model the external toroidal magnetic field is created by the central linear current.
+To calculate the magnetic field created by the thermo-diffusion $\psi$ - Hall current component, we use the way similar to the one, which used for calculation of toroidal field created by the current through the wound up wire.
+If $I$ is the current along the wire, and $n$ is the number of wire small circles on the unit length of the torus big circle, than the external circular magnetic field $B_{\varphi}(r)$ inside the torus, at $R_{t}\ll R$, is approximately written as [36]
+$B_{\varphi}^{in}(r)=\frac{4\pi}{c}nI\frac{r}{R},\quad r\approx R.$ (60)
+In the plasma torus there is a circular Hall current density $j_{in,\psi}$, defined in (59).
 The magnetic field formed by the Hall current could be considered as sum of thin circular currents, similar to the wire current in the laboratory torus, considered above.
-The total circular current on the radial distance $r_{0}$ from the center is equal to ($j_{ex,\psi}2\pi r_{0}d\rho$).
-The magnetic field circulation from this current along the torus at the radius $r<r_{0}$ is equal to $2\pi rB$.
-So the magnetic field $dB_{\phi}$ produced by this sheet of current at the radius $r$ is written as:
-$dB_{\phi}(r)=j_{ex,\psi}\frac{r_{0}}{rc}d\rho,\quad R_{in}<r<r_{0}.$ (45)
-It follows from (44), after substituting ($nI$) by ($j_{ex,\psi}(\rho)d\rho$), and $R$ by $r_{0}$.
-The Hall magnetic field from the whole current in the plasma torus is obtained by integration of (45) over the torus thickness as
-$B_{\phi}(r)=\int_{0}^{R-R_{in}}j_{ex,\psi}\frac{r_{0}}{rc}d\rho,=\frac{1}{rc}\int_{0}^{R-R_{in}}j_{ex,\psi}r_{0}d\rho,\qquad R_{in}<r<R.$ (46)
-On the symmetry plane $z=0$ there is a connection $\rho=R-r_{0}$, $d\rho=-dr_{0}$. For constant current density we have after integration:
-[…]
-It is easy to show that the magnetic field produced by the Hall current has a sign opposite to the field from the central current (Fig.2).
-Field decreasing is minimal on the central line, and its maximal value is in the surface of the torus.
-When the external radial electric field $E_{\rho,ex}$ is present in the torus instead of the temperature gradient, then instead of expression in the Eq.(43), we should use the following one:
-$j_{ex,\rho}=\frac{\sigma_{0}E_{\rho,ex}}{1+(\omega\tau_{e})^{2}},\qquad j_{ex,\psi}=-\sigma_{0}\frac{\omega\tau_{e}E_{\rho,ex}}{1+(\omega\tau_{e})^{2}}.$ (48)
+The $\psi$ - current in the torus on the radial distance $r$ from the center is equal to ($j_{in,\psi}2\pi rd\rho$).
+The magnetic field circulation produced by this current along the large circle of the torus torus, is approximately equal to $2\pi R\,dB_{\varphi}^{in}(r)$.
+So the magnetic field $dB_{\phi}^{in}$ produced by this sheet of current at the radius $r$ is written as:
+$dB_{\varphi}(r)=-\frac{4\pi}{c}j_{in,\psi}\frac{r}{R}d\rho,\quad R_{in}<r<R.$ (61)
+It follows from (60), after substituting ($nI$) by ($j_{ex,\psi}(\rho)d\rho$).
+The Hall magnetic field from the whole current in the plasma torus is obtained by integration of (61) over the torus thickness, as
+$B_{\varphi}(r)=-\frac{4\pi}{c}\int_{R-r}^{R-R_{in}}j_{in,\psi}\frac{r}{R}d\rho,=-\frac{4\pi}{Rc}\int_{R-r}^{R-R_{in}}j_{in,\psi}rd\rho,\qquad R_{in}<r<R.$ (62)
+On the symmetry plane $z=0$ there is a connection $\rho=R-r$, $d\rho=-dr$. For a constant current density we have after integration:
+$B_{\varphi}^{in}(r)=-\frac{4\pi}{Rc}\int_{R-r}^{R-R_{in}}j_{in,\psi}(R-\rho)d\rho=-\frac{2\pi}{Rc}j_{in,\psi}(r^{2}-R_{in}^{2}).\quad R_{in}<r<R.$ (63)
+The magnetic field produced by the Hall current has a sign opposite to the field from the central current (Fig.2).
+The induced magnetic field is zero at the surface of the toroid, and has a maximum (negative) value
+$B_{\varphi}(R)=-\frac{2\pi}{c}\frac{j_{in,\psi}}{rc}(R^{2}-R_{in}^{2}),\qquad\qquad\qquad$ (64)
+on the cental circle of the toroid.
+In numerical simulation, at variable coefficients, we’ll use definitions, for the external field of positive sign (55) as $B_{\varphi}(r)=B_{\varphi}^{ex}(r)\approx B_{0}$, and for the internal field (63) as $B_{\varphi}^{in}(r)=B_{1}(r)$.
+We consider a case with a constant heat flux $Q$, through the torus, along its smaller radius $\rho$, and the heat flux through the unit length of the torus $Q_{\rho}$, which is defined as
+$Q_{\rho}=-2\pi\rho\frac{\kappa_{0}}{1+\omega^{2}\tau_{e}^{2}}\frac{dT}{d\rho}.\qquad\qquad$ (65)
+Here the cyclotron frequency $\omega=\frac{eB}{m_{e}c}$ contains $B=B_{0}+B_{1}$.
+Equations (58), (65) cannot be extended until the axis with $\rho=0$, because of singularities at zero radius.
+We presume that the source of a heat is situated along a thin circle along the axis of the torus, and is represented by a uniformly heated wire with a radius $\rho_{1}$, that is much smaller that the small circle radius of torus $\rho_{1}<<R_{t}$.
+The magnetic field induced by the Hall current is similar to solenoid, having a zero magnetic field at the boundary [30].
+The equations governing the distribution of the induced magnetic field $B_{1}$ and temperature $T$ along the $\rho$ coordinate of the torus, are written, in non-dimensional variables, for the case $R_{t}\ll R$, similar to the cylinder in [20], as
+$\frac{db_{1}}{dx}=\frac{3C_{1}T^{1/2}eQ_{\rho}\omega_{B0}(1+b_{1})}{10kcB_{0}x},$
+$\frac{dT}{dx}=-\frac{1+C_{1}^{2}T^{3}\omega_{B0}^{2}(1+b_{1})^{2}}{2\pi xC_{2}T^{5/2}}Q_{\rho}.$ (66)
+This system is solved at given parameter $Q_{\rho}$.
+In non-dimensional parameters with a radius $x=\rho/R_{t}$, the Hall component $b_{1}=B_{1}/B_{0}$, and temperature $\tilde{T}=T/T_{0}$, the cyclotron frequency is written here as $\omega_{B}={eB_{\phi}}/{m_{e}c}={e(B_{0}+B_{1})}/{m_{e}c}=\omega_{B0}(1+b_{1})$.
+We have used also, that ${\lambda_{0}}/{\kappa_{0}}={3e}/{20kT}$,
+The constants $C_{1}$, $C_{2}$ are determined from relations:
+$\tau_{e}=\frac{3(kT)^{3/2}}{4Z^{2}e^{4}n_{N}\Lambda}\sqrt{\frac{m_{e}}{2\pi}}=C_{1}T^{3/2},$ (67)
+$\kappa_{0}=\frac{40\sqrt{2}kn_{e}}{\pi^{3/2}\Lambda n_{N}}\left(\frac{kT}{e^{2}Z}\right)^{2}\left(\frac{kT}{m_{e}}\right)^{1/2}=C_{2}T^{5/2},$ (68)
+$\omega_{B}\tau_{e}=C_{1}T^{3/2}\omega_{B0}(1+b_{1}).$ (69)
+The dimensionl —ess parameters are introduced as
+$N=\frac{3eQ_{\rho}\omega_{B0}C_{1}}{10kcB_{0}}T_{0}^{1/2},\,\,G=\ C_{1}^{2}T_{0}^{3}\omega_{B0}^{2},\,\,E=\frac{2\pi C_{2}T_{0}^{7/2}}{Q_{\rho}}.$ (70)
+Equations (4.4) have following form with new parameters:
+$\frac{db_{1}}{dx}=N\frac{(1+b_{1}){\tilde{T}}^{1/2}}{x},\qquad\frac{d\tilde{T}}{dx}=-\frac{1+G(1+b_{1})^{2}{\tilde{T}^{3}}}{xE{\tilde{T}^{5/2}}}.$ (71)
+We solve equations (71) numerically in the interval ${\rho_{1}}/{R_{t}}=x_{1}\leq x\leq 1$ at boundary conditions:
+$b_{1}(1)=0,\qquad\tilde{T}(x_{1})=1,\qquad x_{1}=10^{-4}.$ (72)
+Results of the solution are presented on the figures (3-8) for the case of plasma parameters in the laboratory facilities.
+Figure 3: Magnetic field in torus small circle induced by the Hall current, for $G=1.2\cdot 10^{-5}$, $E=0.1$, and three variants: $N_{1}=0.8$; $N_{2}=8.5$; $N_{3}=85.2$.
+These values are related to $Z=1$ and include combinations
+$B_{0}=5\cdot 10^{3}$ G, $T_{0}=2\cdot 10^{5}$ K, $\rho_{0}=10^{-4}$ g/cm3 for $N_{1}$;
+$B_{0}=5\cdot 10^{2}$ G, $T_{0}=2\cdot 10^{5}$ K, $\rho_{0}=10^{-5}$ g/cm3 for $N_{2}$;
+$B_{0}=50$ G, $T_{0}=2\cdot 10^{5}$ K, $\rho_{0}=10^{-6}$ g/cm3 for $N_{3}$.
+Figure 4: Temperature distribution in the torus small circle for the same parameters as in Fig. 3
+Figure 5: Magnetic field in the torus small circle, induced by the Hall current, for $E=0.1$ and three variants: $G_{1}=1.2\cdot 10^{-11}$, $N_{1}=0.085$; $G_{2}=1.2\cdot 10^{-9}$, $N_{2}=0.8$; $G_{3}=1.2\cdot 10^{-7}$, $N_{3}=8.5$.
+These values are related to $Z=1$, and include variants $T_{0}=2\cdot 10^{5}$ K, $B_{0}=50$ G ,
+$\rho_{0}=10^{-3}$ g/cm3 for $N_{1}$$G_{1}$;   $\rho_{0}=10^{-4}$ g/cm3 for $N_{2}$$G_{2}$;   $\rho_{0}=10^{-5}$ g/cm3 for $N_{3}$ $G_{3}$.
+Figure 6: Temperature distribution in the torus small circle for the same parameters as in Fig. 5
+Figure 7: Magnetic field in the torus small circle, induced by the Hall current, N=0.8,E=0.1N=0.8,\,\,E=0.1, and three variants: $G_{1}=4.7\cdot 10^{-5}$; $G_{2}=4.7\cdot 10^{-7}$; $G_{3}=4.7\cdot 10^{-9}$.
+These values are related to $Z=1$, and include variants $\rho=10^{-4}$ g/cm3, $T_{0}=2\cdot 10^{5}\,K$,
+$B_{0}=10^{4}$G, for $G_{1}$;  $B_{0}=10^{3}$ G, for $G_{2}$;   $B_{0}=10^{2}$ G, for $G_{3}$.
+Figure 8: Temperature distribution in the torus small circle for the same parameters as in Fig. 7
 5 Discussion